OpenStudy (anonymous):

A railroad tunnel is shaped like a semiellipse as shown below The height of the tunnel at the center is 54 ft and the vertical clearance must be 18 ft at a point 8 ft from the center. Find an equation for the ellipse.

OpenStudy (anonymous):



OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

i know i have the points (0,54) and (8,18)

OpenStudy (anonymous):

\[\frac{x^2}{a^2} + \frac{y^2}{b^2}=1 \]

OpenStudy (anonymous):

we know \(b^2=54\) so we need \(a^2\) plug in the numbers

OpenStudy (anonymous):

\[\frac{8^2}{a^2}+\frac{18^2}{54}=1\] solve for \(a^2\) and that is it

OpenStudy (anonymous):

i got this far, i just dont get how to solve for a.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

would a=24?